OpenStudy (anonymous):

How to find the stationary point of inflection if we know two points on the graph of y= a(x-h)^3 + k: (-4,0) , (0,28) where 'a'(which is the dilation is equal to 1

OpenStudy (anonymous):

I found the two equations k=28+h^3 and k= - (-4-h)^3

OpenStudy (anonymous):

But I don't know how to find k by using both equations

OpenStudy (anonymous):

If the above equations are correct than solving this system of equations is fairly easy. k=28+h^3 k= - (-4-h)^3 Thus 28+h^3= - (-4-h)^3 Now expand the bracket: (-4-h)^3=-64 -3*16h-3*4h^2- h^3 so - (-4-h)^3=h^3+12h^12+48h+64 28+h^3= - (-4-h)^3 becomes 28+h^3=h^3+12h^12+48h+64 Can you continue now?

OpenStudy (anonymous):

Why is it to the power of 12

OpenStudy (anonymous):

That is a misprint, it should be 2

OpenStudy (anonymous):

28+h^3=h^3+12h^2+48h+64

OpenStudy (anonymous):

So basically u use quadratic equation to find the values

OpenStudy (anonymous):

Yep as h^3 cancels

OpenStudy (anonymous):

Thanks andras I appreciate your help bro

OpenStudy (anonymous):

No problem